OpenStudy (anonymous):

A college loan of $28,000 is made at 8% interest, compound annually. After t years, the amount due, A is given by the function A (t) =28,000 (1.08^t) a) After what amount of the time will the amount due reach $42,000? t=_____ years Round to the nearest whole number.

OpenStudy (anonymous):

t=a/p

OpenStudy (campbell_st):

\[42000 = 28000(1.08^t)\] \[42000/28000 = 1.08^t\] 3/2 = 1.08^t take the log of both sides... e or base 10 \[\ln(\frac{3}{2}) = t \times \ln(1.08)\] \[t = \frac{\ln(\frac{3}{2})}{\ln(1.08)}\] solve for t

OpenStudy (anonymous):

Answer is 5